Circular No. 7731

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001et IN MCG –03-51-9 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope. SN 2001et is located at α = 20^{h}19^{m}35^{s}.00, δ = −18^{o}09′49′′.3 (equinox 2000.0), which is 21′′.7 east and 24′′.3 south of the nucleus of MCG –03-51-9. Magnitudes for the new object: Sept. 13.3 UT, [19.0; 23.2, 18.5; Oct. 1.2, 17.8; 9.2, 17.8; 10.2, 17.8.

SUPERNOVA 2001eu IN ANONYMOUS GALAXY G. Miknaitis, Astronomy Department, University of Washington, re-

ports the detection of an apparent supernova (g′ ∼ 19.4, r′ ∼ 20.4) by the Sloan Digital Sky Survey (SDSS) in images taken with the SDSS 2.5m telescope on Sept. 21.9 UT. SN 2001eu, which was also found in images taken on Sept. 19.9 and 21.9, is located at α = 2^{h}37^{m}58^{s}.80, δ = −1^{o}01′40′′.1 (equinox 2000.0), which is 1′′ west and 1′′ south of the nucleus of a galaxy having position end figures 58^{s}.89, 39′′.1. This galaxy is identified in the SDSS Early Data Release at redshift z = 0.135. SN 2001eu is not present in images taken by the SDSS on 1998 Sept. 19.9. Follow-up observations taken at the APO 3.5-m telescope on Oct. 7.7 by A. Miceli, C. Stubbs, R. Covarrubias, and B. Lawton confirm the existence of the supernova.

SUPERNOVA 2001ep IN NGC 1699 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ep (cf. IAUC 7727), obtained by J. Huchra on Oct. 10.48 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3901 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 300 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light.

POSSIBLE NOVA IN CENTAURUS Additional visual magnitude estimates (cf. IAUC 7727): Oct. 5.927

UT, 9.7 (R. Y. Shida, São Paulo, Brazil); 8.541, 10.1 (A. Pearce, Nedlands, W. Australia); 9.509, 10.9 (Pearce); 10.523, 11.2 (Pearce).

2001 October 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green